OpenStudy (anonymous):

If 37.4 grams of water decomposes at 297 Kelvin and 1.30 atmospheres, how many liters of oxygen gas can be produced? Show all of the work used to solve this problem. 2 H2O (l) 2 H2 (g) + O2 (g)

OpenStudy (anonymous):

answer is 19,4 dm3

OpenStudy (anonymous):

Atomic weights: H=1, O=16, H2O=18, O2=32 37.4gH2O x 1molH2O/18gH2O x 1molO2/2molH2O x 22.7LO2/1molO2 = 23.6 L at STP Second step: P1V1T1 = P2V2/T2, so V2 = P1V1T2/T1P2. And then V2 = (1atm)(23.6L)(297K)/(273K)(1.30atm) = 19.7 L to three significant figures

OpenStudy (anonymous):

n(H2O)=m(H2O)/M(H2O)=2,07 mol n(O2)= 1/2 * n(H2O) = 1,035 mol pV=nRT --> V=nRT/p V(O2)= (1,035 mol * 8,314 JK-1mol-1 * 297K)/131722,5 Pa V(O2) = 0,0194 m3 --> 19,4 dm3 or 19,4 liters